One of the first mixtures of metals, called amalgams, used by dentists for tooth fillings consisted of 26.0 grams of silver, 10.8 grams of tin, 2.4 grams of copper, and 0.8 grams of zinc.

How much silver is in a 25.0 gram sample of this analgam?

thank you

Answer by nerdybill(7384)

26.0 grams of silver, 10.8 grams of tin, 2.4 grams of copper, and 0.8 grams of zinc
.
Total weight = 26.0 + 10.8 + 2.4 + 0.8 = 40 grams
.
Of the total weight, 26 grams is made up of silver Or
26/40 * 100 = 65%
.
So, if you have 25 grams of analgam, 65% would be silver:
.65 * 25 = 16.25 grams